OpenStudy (habbababba1):

If f(x) = 4x + 3 and g(x) = 3x, evaluate...

OpenStudy (habbababba1):

If f(x) = 4x + 3 and g(x) = 3x, evaluate...

OpenStudy (habbababba1):

for

OpenStudy (calculusxy):

So we would have \((4x+3) \times 3x\) but now we need to substitute x for 1/2: \(\large (4(\frac{1}{2}) + 3) \times 3(\frac{1}{2})\)

OpenStudy (calculusxy):

Evaluate that.

OpenStudy (calculusxy):

@habbababba1 Were you able to find an answer?